OpenStudy (anonymous):

Given the function f(x)=2x^2+3x-14/x^2-4 find the domain and graph the function

OpenStudy (solomonzelman):

\(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid purple ]{ f(x)~=~\frac{2x^2+3x-14}{x^2-4} }}}\) `LIKE THIS ? `

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \rm f(x) =\frac{(x-2)(2x+7)}{(x-2)(x+2)} }\) \(\Huge\color{blue}{ \rm f(x) =\frac{2x+7}{x+2} }\)

OpenStudy (solomonzelman):

Work it from here. The domain can be everything (all real numbers) except that one number that makes the denominator =0, which is -2

OpenStudy (anonymous):

thanks but idk how to graph it